Hi all, I was hoping to get some help with the following problem with regards to correctly finding an equation to a circle.
The problem involves a wheel rim with the equation:
x^2 + y^2 - 4x + 2y - 436 = 0
However, the wheel rim circle equation (above) does not account for the tyre on the wheel rim. Find the equation that represents the tyre if its sidewall height is 11cm.
Any help and guidance would be great.
Thanks, -Nick.

Answer by scott8148(6628)

the general equation is ___ (x - h)^2 + (y - k)^2 = r^2
this is a circle, centered at (h,k) with a radius of r

completing the squares for x and y ___ x^2 - 4x + 4 + y^2 + 2y + 1 - 436 = 0 + 4 + 1

(x - 2)^2 + (y + 1)^2 = 431

for the tyre ___ r^2 = {[sqrt(431)] + 11}^2
the center point does not change